OpenStudy (anonymous):

2ln(x+1)-1=0. Solve for x.

OpenStudy (anonymous):

\[\ln (x+1)=1/2\]So,\[x+1=\sqrt{e}\]and,\[x=\sqrt{e}-1\]

OpenStudy (anonymous):

2ln(x+1)-1=0 ===>ln(x+1)=1/2 ===> \[\sqrt{e}=x+1\]

OpenStudy (anonymous):

==>\[ x=\sqrt{e} -1\]